# On the Multiary Algebraic Formulation of an Idempotent Symmetric Limit Convex Structure

### v1

@@ Abstract
Keywords: Generalized Mean, Convexity, Convex hull, Duality, Semilattice, Non-Associative Algebraic Structure, Separation Properties.
@@ 1 Introduction
$\mathbb{B}$-convexity was introduced in [14] as a suitable Painlevé-Kuratowski limit of linear convexities. These generalized convexities are derived from a $\varphi_{p}$-transformation defined as $\varphi_{p}(\lambda)=\lambda^{2p+1}$, following the scheme proposed in [6]. When restricted to the Euclidean orthant $\mathbb{R}_{+}^{n}$, $\mathbb{B}$-convexity [^1] The symbol $\mathbb{B}$, used to define a related notion of convexity, may be somewhat confusing given the concept of convexity introduced by Anatole Beck in Banach spaces [5]. Some clarifications regarding the choice of this notation are provided in [14, 13]. More importantly, note that the operation $\boxplus$, pronounced $\mathbb{B}$oxplus,” visually resembles the Chebyshev unit ball (the $\ell_{\infty}$ ball) when depicted in two dimensions using Cartesian coordinates. This ball plays a central role in the proposed framework and arises as the Hausdorff limit of the sequence of $\ell_{p}$ balls as $p\to\infty$. can be expressed in terms of a suitable lattice structure, reducing to Max-Times convexity. Max-Times convexity is homeomorphic to Max-Plus convexity, which arises in tropical mathematics (see references [29], [28], [25], [27]). These convexities involve an idempotent semi-ring algebraic structure. In this context, $\mathbb{B}$-convex functions were analyzed in [1], and Hahn-Banach-type separation properties were established in [17]. Additionally, fixed point results were obtained in [15] and [18], and alternative idempotent convex structures were proposed in [2] and [31].
It is well known, however, that symmetry, associativity, and idempotence are generally incompatible, except for trivial groups that reduce to the neutral element. In [9], a special class of idempotent magmas was considered in which associativity was relaxed to preserve both symmetry and idempotence. This was achieved by introducing a non-associative operation denoted by $\mathbb{B}$ox-plus $\boxplus$, which benefits from an $n$-ary extension, for any $n$. It has been shown that this notion of convexity can be equivalently characterized using the Kuratowski-Painlevé limit of the generalized convex hull of two points, as defined in [14].
@@
With an algebraic definition of $\mathbb{B}$-convexity in the general case, it appears that this also imposes a stronger definition that enjoys many interesting properties. Paralleling some known properties established in the context of usual convexity, we deduce a limit separation theorem from some earlier results established in [10] and [12]. This we do by using the fact that two disjoint $\mathbb{B}$-polytopes can be separated. Proceeding in this direction, we derive a separation result by approximating convex sets by polytopes.
The paper unfolds as follows. Section 2 presents the framework of $\mathbb{B}$-convexity induced by the $\mathbb{B}$ox-plus operator $\boxplus$ and some related algebraic properties. Section 3 first establishes a preliminary inclusion property for $\mathbb{B}$-polytopes using a suitable notion of determinant and some earlier properties noted in [11]. The converse inclusion is shown, and we provide examples demonstrating that the convexities defined in [9] and [14] are not equivalent. However, their algebraic formulations, in any case, involve an idempotent and non-associative algebraic structure. In Section 4, it is shown that compact $\mathbb{B}$-convex sets can be approximated by $\mathbb{B}$-polytopes. Along this line, we derive a separation result in Section 4. We conclude this section by clarifying the external representation of polytopes by a finite number of half-spaces, as considered in [11].
@@ 2.1 Generalized Means and Convexity
In [14], $\mathbb{B}$-convexity is introduced as a limit of linear convexities. More precisely, for all $p\in\mathbb{N}$, a bijection $\varphi_{p}:\mathbb{R}\longrightarrow\mathbb{R}$ defined by:
@@
and $\phi_{p}(x_{1},...,x_{n})=(\varphi_{p}(x_{1}),...,\varphi_{p}(x_{n}))$ was considered; this is basically the approach of Ben-Tal [6] and Avriel [3]. A field structure on $\mathbb{R}$, for which $\varphi_{p}$ becomes a field isomorphism, can then be defined. Given this change of notation via $\varphi_{p}$ and $\phi_{p}$ we can define a $\mathbb{R}$-vector space structure on $\mathbb{R}^{n}$ by: $\lambda\stackrel{{\scriptstyle\varphi_{p}}}{{\cdot}}x=\phi_{p}^{-1}(\varphi_{p}(\lambda)\cdot\phi_{p}(x))=\lambda\cdot x$ for all $\lambda\in\mathbb{R}$ and $x\stackrel{{\scriptstyle\varphi_{p}}}{{+}}y=\phi_{p}^{-1}(\phi_{p}(x)+\phi_{p}(y))$, for all $x,y\in\mathbb{R}^{n}$. We call these two operations the indexed scalar product and the indexed sum (indexed by $\varphi$).
@@
$Co^{\varphi_{p}}(A)=\Big\{\phi_{p}^{-1}\Big(\sum_{j\in[m]}t_{j}^{2p+1}{\cdot}\phi_{p}\big(x^{(j)}\big)\Big):\big(\sum_{j\in[m]}t_{j}^{2p+1}\big)^{\frac{1}{2p+1}}=1,j\in[m],t_{j}\geq 0\Big\}$ (4)
@@
Moreover, for all $L\subset\mathbb{R}$, we simplify the notations denoting $Co^{p}(L)=Co^{\varphi_{p}}(L)$. The Kuratowski-Painlevé [^2] Though this terminology is usual, it seems that Peano also was at the origin of this notion. lower limit of the sequence of sets $\{A^{(p)}\}_{p\in\mathbb{N}}$, denoted $\operatorname*{\mathrm{Li}}_{p\to\infty}A^{(p)}$, is the set of points $x$ for which there exists a sequence $\{x^{(p)}\}$ of points such that $x^{(p)}\in A^{(p)}$ for all $p$ and $x=\lim_{p\to\infty}x^{(p)}$; the Kuratowski-Painlevé upper limit of the sequence of sets $\{A^{(p)}\}_{p\in\mathbb{N}}$, denoted $\operatorname*{\mathrm{Ls}}_{p\to\infty A^{(p)}}$, is the set of points $x$ for which there exists a subsequence $\{x^{(p_{k})}\}_{k\in\mathbb{N}}$ of points such that $x^{(p_{k})}\in A^{(p_{k})}$ for all $k$ and $x=\lim_{k\to\infty}x^{(p_{k})}$ ; a sequence $\{A^{(p_{k})}\}_{k\in\mathbb{N}}$ of subsets of $\mathbb{R}^{n}$ is said to converge, in the Kuratowski-Painlevé sense, to a set $A$ if $\operatorname*{\mathrm{Ls}}_{p\to\infty}A^{(p)}=A=\operatorname*{\mathrm{Li}}_{p\to\infty}A^{(p)}$, in which case we write $A=\operatorname*{\mathrm{Lim}}_{p\to\infty}A^{(p)}$. We now define $\mathbb{B}$-convexity and its characterization in terms of Kuratowski-Painlevé limit of $\varphi_{p}$-convex sets. For a non-empty finite subset $A=\{x^{(1)},...,x^{(m)}\}\subset\mathbb{R}^{n}$ we let $Co^{\infty}(A)=\operatorname*{\mathrm{Ls}}_{p\longrightarrow\infty}Co^{p}(A)$. When $A\subset\mathbb{R}_{+}^{n}$, it was shown in [14] that we have:
@@
where $\vee$ stands for the lattice operation related to the canonical partial order of $\mathbb{R}^{n}.$ Moreover it is also proved that the sequence $\{Co^{p}(A)\}_{p\in\mathbb{N}}$ is Hausdorff convergent. Such a formulation is useful to apprehend Max-Times that is defined on the idempotent semi-module $(\mathbb{R}_{+}^{n},\vee,\cdot)$. This convexity that is homeomorphic to Max-Plus convexity via a logarithmic transformation (see ).
@@
$\mathbb{B}$-convex sets possess a number of remarkable properties. If $L\subset\mathbb{R}_{+}^{n}$ is $\mathbb{B}$-convex (and therefore Max-Times convex), then $L=\lim_{p\to\infty}Co^{p}(L)=\bigcap_{p=0}^{\infty}Co^{p}(L).$ For any set $S\subset\mathbb{R}_{+}^{n}$, we denote $Co^{\infty}(S)=\lim_{p\to\infty}Co^{p}(S).$ Properties of this operation are developed in [14]. In particular, for all $S\subset\mathbb{R}_{+}^{n}$, the set $Co^{\infty}(S)$ is $\mathbb{B}$-convex.
@@ 2.2 An Algebraic Structure Extended to the whole Euclidean Vector Space
In the following it is established that, though the operation $\boxplus$ does not satisfy associativity, it can be extended by constructing a non-associative algebraic structure which returns to a given $n$-tuple a real value. For all $x\in\mathbb{R}^{n}$ and all subset $I$ of [n][n] [^3] For all positive natural numbers $n$, $n]=\{1,...,n\}$., let us consider the map $\xi_{I}[x]:\mathbb{R}\longrightarrow\mathbb{Z}$ defined for all $\alpha\in\mathbb{R}$ by
@@
For all positive natural numbers $n$ and for all subsets $I$ of $[n]$, let $\digamma_{I}:\mathbb{R}^{n}\longrightarrow\mathbb{R}$ be the map defined for all $x\in\mathbb{R}^{n}$ by
@@
It has been equivalently established in [9] that a set $C\subset\mathbb{R}^{n}$ is equivalently idempotent symmetric convex set if for all $x,y\in C$, $Co^{\infty}(\{x,y\})\subset C$. In the remainder we will denote $Co^{\infty}(\{x,y\})=Co^{\infty}(x,y)$. Moreover, for all $x,y\in\mathbb{R}^{n}$, $Co^{\infty}(x,y)$ is the smallest idempotent symmetric convex set that contains $x$ and $y$. In general, for any subset $E$ of $\mathbb{R}^{n}$, we denote by $\mathrm{IS}[E]$ the smallest idempotent, symmetric, and convex set containing $E$.
@@ 2.3 Limit Systems and Determinant
It is easy to see that the map $x\mapsto A\stackrel{{\scriptstyle p}}{{\cdot}}x$ is $\varphi_{p}$-linear. Conversely, if $g$ is a $\varphi_{p}$-linear map, then it can be represented by a matrix $A$ such that $g(x)=A\stackrel{{\scriptstyle p}}{{\cdot}}x$ for all $x\in\mathbb{R}^{n}$. Notice that the identity matrix $I$ is invariant with respect of $\Phi_{p}$.
@@
This implies that is $f$ is a linear endomorphism having a matrix representation $A$, then $|f^{(p)}|_{p}=|A|_{p}$.
Let $f:\mathbb{R}^{n}\longrightarrow\mathbb{R}^{n}$ be a linear endomorphism having a matrix representation $A$. For all $p\in\mathbb{N}$, let $f^{(p)}$ be its $\varphi_{p}$-linear transformation. Then:
@@
Let $f:\mathbb{R}^{n}\longrightarrow\mathbb{R}^{n}$ be a linear endomorphism having a matrix representation $A$. For all $p\in\mathbb{N}$, let $f^{(p)}$ be its $\varphi_{p}$-linear transformation. If $|A|_{\infty}\not=0$, then there is some $p_{0}\in\mathbb{N}$ such that for all $p\geq p_{0}$ $f^{(p)}$ is $\varphi_{p}$-invertible and for all $b\in\mathbb{R}^{n}$, there exists a solution $x^{(p)}$ to the system $A\stackrel{{\scriptstyle p}}{{\cdot}}x=b$ expressed as:
@@ 3 Algebraic formulation
$x=\sum_{j\in[n+1]}t_{i}^{(p)}x^{(j)}=\frac{1}{|\Lambda|_{p}}\stackrel{{\scriptstyle\varphi_{p}}}{{\sum_{j\in[n+1]}}}|\Lambda_{x}^{(i)}|_{p}x^{(j)}=\frac{1}{|\Lambda|_{p}}\stackrel{{\scriptstyle\varphi_{p}}}{{\sum_{j\in[n+1]}}}\stackrel{{\scriptstyle\varphi_{p}}}{{\sum_{\sigma\in\mathfrak{S}_{n+1}}}}\prod_{i\in[n+1]}\mathrm{sgn}(\sigma)\lambda_{i,\sigma(i)}^{(j)}x^{(j)}.$ (27)
@@
In addition, note that since $t_{j}^{(p)}\leq 0$ for all $p$, we have $\operatorname*{\text{\Huge${}_{\boxplus}$}}_{\sigma\in\mathfrak{S}_{n+1}}\alpha_{\sigma,j}\geq 0$. Moreover, since $\stackrel{{\scriptstyle\varphi_{p}}}{{\sum\limits_{j\in[n+1]}}}t_{j}^{(p)}=1$, implies that $\operatorname*{\text{\Huge${}_{\boxplus}$}}_{\begin{subarray}{c}j\in[n+1]\\ \sigma\in\mathfrak{S}_{n+1}\end{subarray}}\alpha_{\sigma,j}=\lim_{p\longrightarrow\infty}\stackrel{{\scriptstyle\varphi_{p}}}{{\sum\limits_{j\in[n+1]}}}t_{j}^{(p)}=1.$ One can recognize the form of the standard convex hull, except that, for each $j$, the $\alpha_{\sigma,j}$ cannot be grouped due to the lack of associativity. However, in this particular situation, we have assumed that, for each $p$, $x\in Co^{p}(A)$; this is not true in general. Moreover, note that selecting, for each $p$, a point $x^{(p)}$ such that $x^{(p)}\to x$ is not sufficient, as there is no guarantee that the generalized sum will converge to the $\mathbb{B}$ox-Plus sum —its components are no longer independent of $p$.

### v2

@@ Abstract
Keywords: Generalized Mean, Convexity, Convex hull, Duality, Semilattice, Multiary and Non-Associative Algebraic Structure, Separation Properties.
@@ 1 Introduction
$\mathbb{B}$-convexity was introduced in [14] as a suitable Painlevé-Kuratowski limit of linear convexities. These generalized convexities are derived from a $\varphi_{p}$-transformation defined as $\varphi_{p}(\lambda)=\lambda^{2p+1}$, following the scheme proposed in [6]. When restricted to the Euclidean orthant $\mathbb{R}_{+}^{n}$, $\mathbb{B}$-convexity [^1] The symbol $\mathbb{B}$, used to define a related notion of convexity, may be somewhat confusing given the concept of convexity introduced by Anatole Beck in Banach spaces [5]. Some clarifications regarding the choice of this notation are provided in [14, 13]. More importantly, note that the operation $\boxplus$, pronounced $\mathbb{B}$oxplus,” visually resembles the Chebyshev unit ball (the $\ell_{\infty}$ ball) when depicted in two dimensions using Cartesian coordinates. This ball plays a central role in the proposed framework and arises as the Hausdorff-Pompeiu limit of the sequence of $\ell_{p}$ balls as $p\to\infty$. can be expressed in terms of a suitable lattice structure, reducing to Max-Times convexity. Max-Times convexity is homeomorphic to Max-Plus convexity, which arises in tropical mathematics (see references [30], [27], [24], [26]). These convexities involve an idempotent semi-ring algebraic structure. In this context, $\mathbb{B}$-convex functions were analyzed in [1], and Hahn-Banach-type separation properties were established in [17]. Additionally, fixed point results were obtained in [15] and [18], and alternative idempotent convex structures were proposed in [2] and [32].
It is well known, however, that symmetry, associativity, and idempotence are generally incompatible, except for trivial groups that reduce to the neutral element. In [9], a special class of idempotent magmas was considered in which associativity was relaxed to preserve both symmetry and idempotence. This was achieved by introducing a non-associative operation denoted by $\mathbb{B}$ox-plus $\boxplus$, which benefits from an $n$-ary extension, for any $n$. It has been shown that this notion of convexity can be equivalently characterized using the Painlevé-Kuratowski limit of the generalized convex hull of two points, as defined in [14].
@@
With an algebraic definition of $\mathbb{B}$-convexity in the general case, it appears that this also imposes a stronger definition that enjoys many interesting properties. Paralleling some known properties established in the context of usual convexity, we deduce a limit separation theorem from some earlier results established in [10] and [12]. To achieve this objective, we rely on the fact that the separation of two $\mathbb{B}$-polytopes follows directly from the standard separation theorem for polytopes, upon taking a suitable limit. Proceeding in this direction, we derive a separation result by approximating convex sets by polytopes.
The paper unfolds as follows. Section 2 presents the framework of $\mathbb{B}$-convexity induced by the $\mathbb{B}$ox-plus operator $\boxplus$ and some related algebraic properties. Section 3 first establishes a preliminary inclusion property for $\mathbb{B}$-polytopes using a suitable notion of determinant and some earlier properties noted in [11]. Then, the converse inclusion is shown. We deduce a multiary algebraic form of $\mathbb{B}$-convexity that involves an idempotent, non-associative algebraic structure, extending the formalism proposed in [9] to an arbitrary number of points. Moreover, we provide examples demonstrating that that these limiting polytopes do not satisfy the idempotent symmetrical convex structure defined in [9]. In Section 4, it is shown that compact $\mathbb{B}$-convex sets can be approximated by $\mathbb{B}$-polytopes. Along this line, we derive a separation result in Section 4. We conclude this section by clarifying the external representation of polytopes by a finite number of half-spaces, as considered in [11].
@@ 2.1 Generalized Means and Convexity
In [14], $\mathbb{B}$-convexity is introduced as a limit of linear convexities. More precisely, for any $p\in\mathbb{N}$, a bijection $\varphi_{p}:\mathbb{R}\longrightarrow\mathbb{R}$ defined by:
@@
and $\phi_{p}(x_{1},...,x_{n})=(\varphi_{p}(x_{1}),...,\varphi_{p}(x_{n}))$ was considered; this is basically the approach of Ben-Tal [6] and Avriel [3]. A field structure on $\mathbb{R}$, for which $\varphi_{p}$ becomes a field isomorphism, can then be defined. Given this change of notation via $\varphi_{p}$ and $\phi_{p}$ we can define an $\mathbb{R}$-vector space structure on $\mathbb{R}^{n}$ by: $\lambda\stackrel{{\scriptstyle\varphi_{p}}}{{\cdot}}x=\phi_{p}^{-1}(\varphi_{p}(\lambda)\cdot\phi_{p}(x))=\lambda\cdot x$ for all $\lambda\in\mathbb{R}$ and $x\stackrel{{\scriptstyle\varphi_{p}}}{{+}}y=\phi_{p}^{-1}(\phi_{p}(x)+\phi_{p}(y))$, for all $x,y\in\mathbb{R}^{n}$. We call these two operations the indexed external product and the indexed sum (indexed by $\varphi_{p}$).
@@
$Co^{\varphi_{p}}(A)=\Big\{\phi_{p}^{-1}\Big(\sum_{j\in[m]}t_{j}^{2p+1}{\cdot}\phi_{p}\big(x^{(j)}\big)\Big):\big(\sum_{j\in[m]}t_{j}^{2p+1}\big)^{\frac{1}{2p+1}}=1,j\in[m],t_{j}\geq 0\Big\}.$ (4)
@@
Moreover, for all $L\subset\mathbb{R}$, we simplify the notations denoting $Co^{p}(L)=Co^{\varphi_{p}}(L)$. The Kuratowski-Painlevé [^2] We have adopted the standard terminology “Painlevé–Kuratowski limits” throughout the manuscript. We also note that the terminology “Peano–Kuratowski” appears in the historical literature; see, for instance, the comments in [28, Chapter 4] and in [31, Volume I, Chapter 1, Comments], where these terminological issues are discussed. lower limit of the sequence of sets $\{A^{(p)}\}_{p\in\mathbb{N}}$, denoted $\operatorname*{\mathrm{Li}}_{p\to\infty}A^{(p)}$, is the set of points $x$ for which there exists a sequence $\{x^{(p)}\}_{p\in\mathbb{N}}$ of points such that $x^{(p)}\in A^{(p)}$ for all $p$ and $x=\lim_{p\to\infty}x^{(p)}$; the Kuratowski-Painlevé upper limit of the sequence of sets $\{A^{(p)}\}_{p\in\mathbb{N}}$, denoted $\operatorname*{\mathrm{Ls}}_{p\to\infty}A^{(p)}$, is the set of points $x$ for which there exists a subsequence $\{x^{(p_{k})}\}_{k\in\mathbb{N}}$ of points such that $x^{(p_{k})}\in A^{(p_{k})}$ for all $k$ and $x=\lim_{k\to\infty}x^{(p_{k})}$ ; a sequence $\{A^{(p_{k})}\}_{k\in\mathbb{N}}$ of subsets of $\mathbb{R}^{n}$ is said to converge, in the Kuratowski-Painlevé sense, to a set $A$ if $\operatorname*{\mathrm{Ls}}_{p\to\infty}A^{(p)}=A=\operatorname*{\mathrm{Li}}_{p\to\infty}A^{(p)}$, in which case we write $A=\operatorname*{\mathrm{Lim}}_{p\to\infty}A^{(p)}$. We now define $\mathbb{B}$-convexity and its characterization in terms of Kuratowski-Painlevé limit of $\varphi_{p}$-convex sets. For a non-empty finite subset $A=\{x^{(1)},...,x^{(m)}\}\subset\mathbb{R}^{n}$ we let $Co^{\infty}(A)=\operatorname*{\mathrm{Ls}}_{p\longrightarrow\infty}Co^{p}(A)$. When $A\subset\mathbb{R}_{+}^{n}$, it was shown in [14] that we have [^3] In the following, we denote $\mathbb{R}_{+}=[0,+\infty[$ and $\mathbb{R}_{++}=]0,+\infty[$.:
@@
where $\vee$ stands for the lattice operation related to the canonical partial order of $\mathbb{R}^{n}.$ Moreover it is also proved that the sequence $\{Co^{p}(A)\}_{p\in\mathbb{N}}$ is Hausdorff-Pompeiu convergent (see [28]and [31] for definitions and terminologies). Such a formulation is useful to apprehend Max-Times convexity that is defined on the idempotent semi-module $(\mathbb{R}_{+}^{n},\vee,\cdot)$. This convexity is homeomorphic to Max-Plus convexity via a logarithmic transformation (see [15] ).
@@
$\mathbb{B}$-convex sets possess a number of remarkable properties. If $L\subset\mathbb{R}_{+}^{n}$ is $\mathbb{B}$-convex (and therefore Max-Times convex), then $L=\lim_{p\to\infty}Co^{p}(L)=\bigcap_{p=0}^{\infty}Co^{p}(L).$ For any set $S\subset\mathbb{R}_{+}^{n}$, we denote $Co^{\infty}(S)=\operatorname*{\mathrm{Lim}}_{p\to\infty}Co^{p}(S).$ Properties of this operation are developed in [14]. In particular, for all $S\subset\mathbb{R}_{+}^{n}$, the set $Co^{\infty}(S)$ is $\mathbb{B}$-convex.
@@ 2.2 An Algebraic Structure Extended to the whole Euclidean Vector Space
In the following it is established that, though the operation $\boxplus$ does not satisfy associativity, it can be extended by constructing a non-associative algebraic structure which returns to a given $n$-tuple a real value. For all $x\in\mathbb{R}^{n}$ and any subset $I$ of $[n]$, let us consider the map $\xi_{I}[x]:\mathbb{R}\longrightarrow\mathbb{Z}$ defined for all $\alpha\in\mathbb{R}$ by
@@
For all positive natural numbers $n$ and for any subset $I$ of $[n]$, let $\digamma_{I}:\mathbb{R}^{n}\longrightarrow\mathbb{R}$ be the map defined for all $x\in\mathbb{R}^{n}$ by
@@
It has been equivalently established in [9] that a set $C\subset\mathbb{R}^{n}$ is equivalently idempotent symmetric convex set if for all $x,y\in C$, $Co^{\infty}(\{x,y\})\subset C$. In the remainder we will denote $Co^{\infty}(\{x,y\})=Co^{\infty}(x,y)$. Moreover, for all $x,y\in\mathbb{R}^{n}$, $Co^{\infty}(x,y)$ is the smallest idempotent symmetric convex set that contains $x$ and $y$. In general, for any subset $E$ of $\mathbb{R}^{n}$, we denote by $\mathrm{IS}[E]$ the smallest idempotent symmetric convex set containing $E$.
@@ 2.3 Limit Systems and Determinant
If $f:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is a linear endomorphism, then the map $f^{(p)}:\mathbb{R}^{n}\to\mathbb{R}^{n}$, defined for all $x\in\mathbb{R}^{n}$ by $f^{(p)}(x)=\phi_{p}^{-1}\circ f\circ\phi_{p}(x)$, is called the $\varphi_{p}$-linear transformation associated with $f$. By construction, this map is $\varphi_{p}$-linear. It is easy to see that the map $x\mapsto A\stackrel{{\scriptstyle p}}{{\cdot}}x$ is $\varphi_{p}$-linear. Conversely, if $g$ is a $\varphi_{p}$-linear map, then it can be represented by a matrix $A$ such that $g(x)=A\stackrel{{\scriptstyle p}}{{\cdot}}x$ for all $x\in\mathbb{R}^{n}$. Notice that the identity matrix $I$ is invariant with respect of $\Phi_{p}$.
@@
This implies that if $f$ is a linear endomorphism having a matrix representation $A$, then $|f^{(p)}|_{p}=|A|_{p}$. Moreover, we have:
@@
Let $f:\mathbb{R}^{n}\longrightarrow\mathbb{R}^{n}$ be a linear endomorphism having a matrix representation $A$. For all $p\in\mathbb{N}$, let $f^{(p)}$ be its $\varphi_{p}$-linear transformation. If $|A|_{\infty}\not=0$, then there is some $p_{0}\in\mathbb{N}$ such that, for all $p\geq p_{0}$, $f^{(p)}$ is $\varphi_{p}$-invertible and for all $b\in\mathbb{R}^{n}$, there exists a solution $x^{(p)}$ to the system $A\stackrel{{\scriptstyle p}}{{\cdot}}x=b$ expressed as:
@@ 3 Algebraic formulation
$x=\sum_{j\in[n+1]}t_{i}^{(p)}x^{(j)}=\frac{1}{|\Lambda|_{p}}\stackrel{{\scriptstyle\varphi_{p}}}{{\sum_{j\in[n+1]}}}|\Lambda^{(j)}|_{p}x^{(j)}=\frac{1}{|\Lambda|_{p}}\stackrel{{\scriptstyle\varphi_{p}}}{{\sum_{j\in[n+1]}}}\stackrel{{\scriptstyle\varphi_{p}}}{{\sum_{\sigma\in\mathfrak{S}_{n+1}}}}\prod_{i\in[n+1]}\mathrm{sgn}(\sigma)\lambda_{i,\sigma(i)}^{(j)}x^{(j)}.$ (27)
@@
In addition, note that since $t_{j}^{(p)}\geq 0$ for all $p$, we have $\operatorname*{\text{\Huge${}_{\boxplus}$}}_{\sigma\in\mathfrak{S}_{n+1}}\alpha_{\sigma,j}\geq 0$. Moreover, since $\stackrel{{\scriptstyle\varphi_{p}}}{{\sum\limits_{j\in[n+1]}}}t_{j}^{(p)}=1$, this implies that $\operatorname*{\text{\Huge${}_{\boxplus}$}}_{\begin{subarray}{c}j\in[n+1]\\ \sigma\in\mathfrak{S}_{n+1}\end{subarray}}\alpha_{\sigma,j}=\lim_{p\longrightarrow\infty}\stackrel{{\scriptstyle\varphi_{p}}}{{\sum\limits_{j\in[n+1]}}}t_{j}^{(p)}=1.$ One can recognize the form of the standard convex hull, except that, for each $j$, the $\alpha_{\sigma,j}$ cannot be grouped due to the lack of associativity. However, in this particular situation, we have assumed that, for each $p$, $x\in Co^{p}(A)$; this is not true in general. Moreover, note that selecting, for each $p$, a point $x^{(p)}$ such that $x^{(p)}\to x$ is not sufficient, as there is no guarantee that the generalized sum will converge to the $\mathbb{B}$ox-Plus sum —its components are no longer independent of $p$.